40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14105     Entry time: Thu Jul 26 01:52:01 2018
Author: terra 
Type: Update 
Category: Thermal Compensation 
Subject: heater work update 

Just a quick update: over the past few days we've taken (at least) 5 scans around each peak [carrier - HOM3] at 9.4V/0.8A, 4 scans around [carrier - HOM5] at 12V/0.9A hot state with the reflector setup. We also have (at least) 5 scans of carrier - HOM5 in cold state. I attach a rough overview of the peak magnitude shifts in the first attachment. Analysis ongoing. All data stored in annalisa/postVent/{date}

Initial shifts just based on rought peak placement in the meantime:

            [9.4V/0.8A]   [12V/0.9A]

HOM1    10 kHz         20 kHz

HOM2    18 kHz         28 kHz

HOM3     30 kHz        40 kHz

HOM4     N/A             26 kHz

HOM5     N/A             35 kHz

I also attach the heating thermal transient from today (12V/0.9A) as seen by the opLevs. We see a shorter time constant for pitch, longer for yaw, preceeded by a dip in yaw. Similar behavior yesterday for slightly less heating, though less pronounced pre-dip. The heater is offcentered on the optic horizontally; likely this is part of the induced yaw. The spikey stuff i removed is from people walking around inside during the transient.

I've left the heater and LSC off for the night. Heater off at 2:07 am local time.

Please don't touch the oplevs; we're taking a cool down measurement.

Attachment 1: OpLev_thermal_drift.pdf  646 kB  Uploaded Thu Jul 26 03:01:01 2018  | Hide | Hide all | Show all
OpLev_thermal_drift.pdf
Attachment 2: hotColdAll.pdf  60 kB  Uploaded Thu Jul 26 03:19:30 2018  | Show | Hide all | Show all
ELOG V3.1.3-